OpenStudy (anonymous):

The length of each side of a square is 5 inches more than the length of each side of a smaller square. The difference of the areas of the squares is 95 inches squared. Find the lengths of the sides of the two squares.

OpenStudy (anonymous):

x^2 + 10x + 25 - x^2 = 95 x^2 - x^2 + 10x = 95 - 25 10x = 70 x = 7 inches is the side of the smaller square then 7 + 5 = 12 inches is the side of the larger square : : Check this by finding the areas 12^2 - 7^2 144 - 49 = 95; confirms our solutions

OpenStudy (anonymous):

where did 10x and 25 come from?

OpenStudy (anonymous):

if the smaller square has sides of length x, then the larger square has sides of length x + 5. Area of the smaller square is x^2. area of the larger square is (x + 5)(x + 5), or x^2 + 10x + 25.

OpenStudy (anonymous):

ahhhh okay, thanks.

OpenStudy (anonymous):

ohsnapigotamedal

OpenStudy (anonymous):

lol you were helpful :D